OpenStudy (anonymous):

use sinx=1/4 to evaluate the following function. (a) sin(-x)= (b) csc(-x)=

OpenStudy (jdoe0001):

sin(-x) = - sin(x) \(\bf csc(x) = \cfrac{1}{sin(x)}\implies csc(-x) = \cfrac{1}{-\quad sin(x)} \)

OpenStudy (ikram002p):

(a) \[\sin -x =\sin \pi -\sin x = 0- 1/4 =-1/4\] (b) \[\csc -x =\frac{ 1 }{ \sin -x } = \frac{ 1 }{ \frac{- 1 }{ 4} } = -4 \]